Question 1


a) What is the concentration of Ca2+ in a saturated solution of CaF2 in pure water, given that Ksp = 4.0 x10^-11?

Write the chemical equation for this reaction.


b) What is the concentration of Ca2+ in a saturated solution of CaF2 in 0.361 M NaF?

Write the chemical equation. SHow your use of the ICE table.


c) Without doing any calculations, would you expect the solution to be in part b) to be more acidic, basic, or neutral? Explain why it would be acidic, basic, or neutral by writing the chemical equation.


1
Expert's answer
2022-03-13T15:21:29-0400

(a) The equilibrium reaction can be given as follows

"CaF_2 \longleftrightarrow Ca^{2+} + 2F^{-1}"


"K_{sp}= [Ca^{2+}][F^{-1}]^2"


"4.01\u00d710^{-11}=S(2S)^2"


S= 3.17 ×10-6 M

(b)

"[F^{-1}]=0.361+x"

"[Ca^{2+}]=x"

"K_{sp}= [Ca^{2+}][F^{-1}]^2"

"4.01\u00d710^{-11}= [x][.361+x]^2"

x= 2.9×10-10 M

(c)

By looking concentration of hydrogen ions so produced we can identified whether a solution is acidic or basic